Constants Vector A has components A: 1.40 cm , A,- 2.35 cm : vector B has components B 4.15 cm B,-3.76 cm

media%2F59a%2F59a9efb9-f7ad-4fdb-ad79-a5

media%2F29b%2F29bd00ab-a5cd-4b1d-ad6e-99

media%2F035%2F0350909f-ee23-46d6-a615-f4

media%2Fcdd%2Fcddee5c8-9e24-425e-8ecc-28

media%2F19b%2F19b9dcaa-e12d-4d2e-91c5-33



media%2F397%2F3971382e-d284-4e13-8c35-39
Answer #1

Given,

A = 1.40 i + 2.35 j

B = 4.15 i - 3.76 j

A)

A + B = (5.55 i - 1.41 j) cm

B)

Magnitude, A + B = sqrt (5.55^2 + 1.41^2) = 5.73 cm

C)

Angle, theta = arctan (-1.41/5.55) = - 14.25 degree

Counter Clockwise angle, theta = 360 - 14.25 = 345.75 degree

D)

Vector, B - A = (2.75 i - 6.11 j) cm

E)

Magnitude, B - A = sqrt(2.75^2 + 6.11^2) = 6.70 cm

F)

Angle, theta = arctan (-6.11/2.75) = - 65.77 degree

Counter Clockwise angle, theta = 360 - 65.77 = 294.23 degree

Comment in case any doubt please rate my answer ....